# Divide et Impera: Learning impedance families for peg-in-hole assembly

### v1

@@ II-F1 Neural Networks for Binary Classification
To investigate Claim 4—that common assembly strategies applied to all pegs—we used a neural network for binary classification. Our goal was to find the best-fit model that predicted success for our experimental data with $N=4*3^{8}$ data points and labels $y=\{0,1\}$, i.e., true (1) and false (2). The “best-fit” was determined by the minimum of the Binary Cross Entropy:
@@
We used a binary cost, as only success or failure is important for industrial peg-in-hole assembly, not the minimization of a continuous error function such as the Euclidean error.
For given input impedance parameters $\bm{x}=\{k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}\}$, we wanted to find the function $f(\bm{x})$ that represents the boundary of our data, dividing it into success or failure. For the neural network, we chose the following structure with two hidden layers:
The network was trained for all pegs, i.e., the eight impedance values and their success flags were merged into one dataset. Once the network was trained, we used the network to produce successful data and analyze whether all pegs shared a common assembly strategy, represented by a lower-dimensional PC space (sec. III-D).
The two hidden layers and output of the neural network were defined as follows: the first hidden layer transformed the eight-dimensional normalized input features $\mathbf{X}^{(i)}=\{k_{x}^{(i)},k_{y}^{(i)},k_{z}^{(i)},k_{C}^{(i)},k_{B}^{(i)},k_{A}^{(i)},\zeta_{t}^{(i)},\zeta_{r}^{(i)}\}\in\mathbb{R}_{>0}^{8}$ to 64 neurons; the second hidden layer took the 64 features from the hidden layer and passed them to another 64 neurons; the output layer transformed the 64 outputs of the second layer to the predicted output $\hat{y}^{(i)}$. After each of the first two layers, a Rectified Linear Unit (ReLU) activation function was used to introduce non-linearity into the model. The Sigmoid activation function was used at the output layer to map the final value to a probability between 0 and 1. The Adam optimizer [48] was used to update the weights of the network with a fixed learning rate of $1\times 10^{-4}$.
@@ III Results
III-A Claim 1: Success with a range of parameters
@@ III-A1 Success Rate
Tab. I reports the success rates for every peg as well as the absolute number of total successes. Excluding the square, which presented the highest success rate ($9.9\%$), the other three pegs presented a success rate ranging between $1.5\%$ and $2.8\%$. We hypothesize that the success rate of the square peg was substantially higher because we used this peg for the initial programming of the application, which determined the kinematic primitives (translations, rotations, and oscillation) and were unchanged for the other pegs (sec. II). In an absolute sense, among the $6561$ tested impedance parameters, it was possible to find between $101$ (triangle) and $646$ (square) feasible peg insertion solutions.
This first result confirms that—given a certain assembly strategy common across all pegs—multiple impedance solutions exist for the same peg. This finding suggests that there is a vast space of ”good-enough“ solutions for a given task, which contrasts with optimization methods that produce only a single set of impedance parameters [17, 49].
@@ III-A2 Success Rate Distribution
For each of the pegs, the distribution of success rates for each impedance parameter ($k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}$) was computed by dividing each impedance parameter into a number of equally spaced bins and computing the success rate for each bin. Fig. 3 shows the observed distributions. A linear model was fit to each success rate distribution in each of the impedance parameters and for each peg. The coefficient of determination $R^{2}$ was computed, and the resulting linear fit was statistically tested against the degenerate null slope.
Interestingly, different pegs shared similar success rate distributions. The success rate grew with increase in vertical stiffness $k_{z}$ in all pegs, and the linear models presented good ($R^{2}>0.6$) coefficients of determination ($R^{2}_{triangle}=0.86,R^{2}_{square}=0.91,R^{2}_{hexagon}=0.77,R^{2}_{cylinder}=0.65$) with significant slopes ($p<0.05$) for all pegs. A similar behavior was observed for the success rate and the $k_{C}$ rotational stiffness across all pegs. The linear models presented high coefficients of determination ($R^{2}_{triangle}=0.73,R^{2}_{square}=0.85,R^{2}_{hexagon}=0.94,R^{2}_{cylinder}=0.88$) and significant slopes ($p<0.05$) for all pegs.
Another interesting finding was that the success rate appeared to be independent of the damping ratio parameters ($\zeta_{t},\zeta_{r}$). In fact, linear fits of the success rate to these two parameters presented low coefficients of determination and non-significant slopes ($p>0.05$) for all but two cases ($\zeta_{r_{square}},\zeta_{t_{hexagon}}$). Tab. II presents the complete details of the linear model results.
The success rate for the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$) was less consistent across pegs. Noticeably, the success rate distributions over $k_{x},k_{y}$ in the triangle peg and the cylinder peg were similar, with a concentration of success for low values of these two impedance parameters. However, the same was not true for the square and hexagon pegs. Appendix D contains the complete details of the linear model fitting performance on the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$).
@@
Figure 3: Distribution of success rates (blue bars) for each impedance parameter fpr the four different pegs (triangle, square, hexagon, and cylinder). For each histogram, the red solid lines represent the best linear fit, while the red dashed lines show the confidence intervals. An asterisk (*) was added for those linear fits that presented a significant slope.

### v2

@@ II-E4 Neural network to reveal common assembly strategies
To investigate common assembly strategies applied to all pegs, we used a neural network for binary classification. Our goal was to find the best-fit model that predicted success for our experimental data with $N=4\cdot 3^{8}$ data points and labels $y=\{0,1\}$, i.e., true (1) and false (2). The “best-fit” was determined by the minimum of the Binary Cross Entropy:
@@
We used a binary cost, as only success or failure is important for assembly, not the minimization of a continuous error function such as the Euclidean error.
For given input impedance parameters $\bm{x}=\{k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}\}$, we wanted to find the function $f(\bm{x})$ that represents the boundary of our data, dividing it into success or failure.
The network was trained for all pegs, i.e., the eight impedance values and their success flags were merged into one dataset. Once the network was trained, we used the network to produce successful data and analyze whether all pegs shared a common assembly strategy, represented by a lower-dimensional PC space (Sec. III-D). More details about the neural network structure can be found in Appendix F.
@@ III Results
This section presents the findings of our investigation into the role of impedance parameters in robotic peg-in-hole assembly. We highlight task-specific and generalized strategies derived through PCA and K-means Clustering. We introduce a neural network model designed to predict and visualize successful impedance parameters.
III-A Success rate and distribution
@@ III-A1 Success Rate
Tab. I presents the success rates for each peg, along with the absolute number of successful insertions. The square peg achieved the highest success rate ($9.9\%$), while the other three pegs exhibited significantly lower success rates, ranging from $1.5\%$ to $2.8\%$. We hypothesize that the square peg’s higher success rate is due to its use in the initial programming phase, where kinematic primitives—such as translations, rotations, and oscillations—were optimized and remained unchanged for the other peg shapes.
The overall low success rates highlight the difficulty of tuning impedance parameters for contact-rich tasks, emphasizing the challenge faced by robot programmers in identifying feasible parameter sets. Despite these challenges, our results show that among the $6,561$ tested impedance parameters, a considerable number of successful solutions were found, ranging from $101$ for the triangular peg to $646$ for the square peg, reinforcing the importance of structured impedance selection strategies in robotic assembly.
Interestingly, despite the complexity of impedance tuning, our findings demonstrate that multiple successful impedance solutions exist for each peg shape, even when a fixed assembly strategy is applied uniformly across all pegs. This suggests that rather than a single optimal impedance value, there is a broader range of viable solutions, complementing optimization methods that typically focus on identifying a single best set of impedance parameters [17, 61].
@@ III-A2 Success Rate Distribution
Figure 4: Distribution of success rates (blue bars) for each impedance parameter for the four different pegs (triangle, square, hexagon, and cylinder). For each histogram, the red solid lines represent the best linear fit, while the red dashed lines show the confidence intervals. An asterisk (*) was added for those linear fits that presented a significant slope.
For each of the pegs, the distribution of success rates for each impedance parameter ($k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}$) was computed by dividing each impedance parameter into a number of equally spaced bins and computing the success rate for each bin.
Fig. 4 shows the observed distributions. Interestingly, different pegs shared similar success rate distributions. The success rate grew with increase in vertical stiffness $k_{z}$ in all pegs, and the linear models presented good coefficients of determination ($R^{2}_{triangle}=0.86,R^{2}_{square}=0.91,R^{2}_{hexagon}=0.77,R^{2}_{cylinder}=0.65$) with significant slopes ($p<0.05$) for all pegs. A similar behavior was observed for the success rate and the $k_{C}$ rotational stiffness across all pegs. The linear models presented high coefficients of determination ($R^{2}_{triangle}=0.73,R^{2}_{square}=0.85,R^{2}_{hexagon}=0.94,R^{2}_{cylinder}=0.88$) and significant slopes ($p<0.05$) for all pegs.
Another interesting finding was that the success rate appeared to be independent of the damping ratio parameters ($\zeta_{t},\zeta_{r}$). In fact, linear fits of the success rate to these two parameters presented low coefficients of determination and non-significant slopes ($p>0.05$) for all but two cases ($\zeta_{r_{square}},\zeta_{t_{hexagon}}$).
The success rate for the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$) was less consistent across pegs. Noticeably, the success rate distributions over $k_{x},k_{y}$ in the triangle peg and the cylinder peg were similar, with a concentration of success for low values of these two impedance parameters. However, the same was not true for the square and hexagon pegs.
Appendix G presents the complete details of the linear model results and fitting performance on the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$).
III-B Low dimensional representation of solutions
To test whether there existed a non-arbitrary lower dimension subspace to represent the solutions, we projected the impedance solutions into a lower dimensional subspace using PCA. We found that 3 PCs could reasonably describe the data, a substantial reduction from the 8 dimensional solution space. Specifically, 3 PCs accounted for 83.8%, 96.6%, 95.9%, 88.7%, and 95.7% of the successful trials of the cylinder, hex, square, triangle, and all peg shapes, respectively.
To show that the reduction of the solution space was not a random artifact, we compared the slope of the variance-accounted-for by the first three principal components to the slope of the variance-accounted-for by a random subset of $n$ trials from the data, where $n$ is the number solutions reported in Tab. I. The amount of variance-accounted-for by each PC in the solution space is shown Appendix B (Fig. 10).
We conducted a pair-wise t-test between the variance-accounted-for slopes of the successful trials and the random trials. We found that the slopes of the successful trials were significantly steeper; that is, the slopes of the successful trials were significantly more negative ($p<0.001$). This confirms that the PCs extracted from the successful trials were not random artifacts; rather, they represent an impedance solution subspace.
Because the data patterns were effectively captured by three PCs, we were able to visualize them in a 3-dimensional space without losing significant information. By projecting the failure trials into this 3-dimensional subspace, we observed that the failures frequently fell outside the identified solution subspace, particularly in the cylinder and triangle trials (Fig. 5). This visualization underscores the existence of a meaningful, non-arbitrary lower-dimensional solution space to represent the impedance solutions. Moreover, analyzing this solution space highlighted the multitude of successful impedance parameter configurations within EDA, illustrating the flexibility and adaptability of this approach. This will be further elaborated in Sec. III-D.
III-C Task-specific assembly strategies for individual pegs
Figure 5: Left Column: The successful (green) and unsuccessful (red) trials of the cylinder, hex, square, and triangle pegs are projected into the solution space obtained using PCA. Note that often the failures fall outside the PCA-reduced solution space. 2D projections of these graphs can be found in  11. Right Column: The coefficients of each of the three Principal Components are shown.
We used K-means Clustering to identify clusters of different successful assembly strategies in our eight-dimensional impedance solution space. Specifically, we used the silhouette coefficient to identify the “optimal” number of clusters. In Fig. 6 we see that the optimal number of clusters was four, with a silhouette coefficient of 0.33. Further, in Fig. 6 we see that often multiple clusters contained different pegs. Only Cluster 1 (red) contained all four pegs; it was the only cluster to include cylinder solutions. Together, these results suggest that similar strategies enabled different tasks (i.e., different pegs).
To verify the distinctiveness of the identified clusters, we conducted a one-way Multivariate Analysis of Variance (MANOVA), using the cluster group as the independent variable with four levels. The results revealed significant differences among the clusters on the combined dependent variables, indicating that the multivariate means of the clusters spanned at least three dimensions (Appendix E-A). These findings validate the clusters identified by K-means Clustering, confirming distinct multivariate mean vectors for each cluster. This demonstrates the existence of task-specific assembly strategies for individual pegs. We identified families of solutions, revealing 2 clusters for the cylinder peg, 4 for the hex peg, 3 for the square peg, and 2 for the triangle peg (Fig. 12). More information about their silhouette coefficients can be found in appendix D-A.
Figure 6: Top Left: The successful parameters of all pegs. Top right: The silhouette coefficient found from K-means Clustering as a function of the number of clusters. An optimal number of clusters is denoted by a red X. Bottom: The successful trials of all pegs were projected into their solution space obtained using PCA. Each color denotes a different cluster identified by the k-means clustering algorithm. In the legend, “Cl” is an abbreviation for cluster. Different marker shapes denote different pegs.
The right column of Fig. 7 shows the different solution strategies that were identified. For example, in the successful trials of the cylinder peg, Cluster 1 contained lower values of $k_{z}$ while Cluster 2 contained higher values of $k_{z}$. The difference between the 4 clusters in the successful hex trials depended on the combination of $k_{x}$ and $k_{y}$: Cluster 1 contained lower values of $k_{x}$ and higher values of $k_{y}$, Cluster 2 contained higher values of $k_{x}$ and lower values of $k_{y}$, Cluster 3 contained higher values of both $k_{x}$ and $k_{y}$, and Cluster 4 contained lower values of both $k_{x}$ and $k_{y}$. The three clusters identified in the successful square trials contained combinations of values in $k_{x}$, $k_{y}$, and $k_{z}$. Finally, the two clusters identified in the successful triangle trials differed in their range of values in $k_{y}$. All in all, this indicates that through the constrained assembly strategy multiple combinations of $k_{x}$ and $k_{y}$ were feasible. This showcases that families of solutions exist which encode different assembly strategies.
To verify the distinctiveness of the identified clusters, we conducted a one-way Multivariate Analysis of Variance (MANOVA). The results confirmed significant differences among clusters for all peg shapes, indicating that the multivariate means of the clusters occupied distinct subspaces. For example, the cylinder and triangle clusters lay in one-dimensional spaces, the square clusters spanned two dimensions, and the hex clusters required three dimensions (appendix E-B). These findings validate the distinctiveness of the solution clusters and support the conclusion that multiple families of solutions encode different assembly strategies.
III-D Generalized assembly strategies across all pegs
As we have shown in our results, a range of successful parameters exist for each peg (Sec. III-A). These parameters might be grouped into different assembly strategies for each individual peg (Sec. III-C). To examine whether a common assembly strategy existed across all pegs, we created a binary classification neural network to sample over the whole space of successful parameters (Sec. II-E).
[…]
We used the model above to produce 100,000 sets of successful impedance parameters; these data will be referred to as the predicted solutions. A PCA was conducted on a combined dataset of the 100,000 predicted solutions and the previously found successful trials. The projection into the first three PCs is shown in Fig. 8 (note that the clusters here are the same as those in Fig. 6). Three dominant PCs were present, accounting for 96.7% of the variance, while subsequent PCs accounted for less than 3% each.
Some interesting findings can be observed. First, the predicted solution space for all pegs (shown in grey) is in fact not a uniform cube, suggesting that there is a subspace—or a family of solutions—common to all pegs. Note that areas in the solution cube outside of this subspace suggest impedance solutions that are not common to all pegs. For example, Cluster 1 (red) contains a solution strategy encompassing all four pegs. The predicted solution space here contains a higher density of predicted solutions in this area (grey) as opposed to other areas (e.g., Cluster 4); this is best seen in the top right of Fig. 8. This intuition is consistent with our results.
Fig. 9 presents the predicted success rates for all peg types, based on the same 100,000 data points. The predicted success rates for all individual pegs can be found in appendix G-A.
Figure 9: Distribution of success rates for parameters predicted to be successful by the neural network. The distributions were generated by sampling impedance parameters uniformly between the minimum and maximum values specified in Sec. II until 100,000 successes were achieved, resulting in a success rate of $4.32\%$.
Although it is challenging to interpret all solutions due to the neural network predicting outcomes across all pegs, some interesting trends can be observed. The success rate was highest for low linear stiffness values $k_{x}$ and $k_{y}$, suggesting that the peg more easily slips into the hole during assembly with these lower parameters, compensating for imperfect alignment. A similar trend to that observed in Fig. 4 is noted: success increases with higher vertical stiffness $k_{z}$ and rotational stiffness $k_{C}$. Once the peg is aligned with the hole, a certain force must be exerted, hence higher values of $k_{z}$ are more successful. Additionally, the original programmed orientation about the x-axis seemed to be similar across all pegs, which may explain the higher success rates for high rotational stiffness $k_{C}$. Interestingly, $k_{A}$ appears to be successful across all parameter ranges. The rotational stiffness $k_{B}$ shows the highest success rates at intermediate values, which might indicate that the final orientation of the peg and hole was not perfectly programmed and varied slightly across pegs. This could explain the optimal balance of maintaining orientation without being overly stiff about the y-axis. The damping parameters $\zeta_{t}$ and $\zeta_{r}$ were successful across all ranges, with a growing success rate for increasing values of $\zeta_{t}$ and a declining success rate for decreasing values of $\zeta_{r}$. Still, we conclude that for our assembly application, the effect of the selected damping ratio was minor, most likely due to the low speed of the task.
III-E Neural network-based impedance success predictor
Even for a given assembly strategy (submovements and oscillations in algorithm 1), determining a set of feasible impedance parameters $\{k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}\}$ is not trivial [43]. To address this challenge, we trained a neural-network-based success predictor (Sec. III-D) that can assist users—especially those working with peg geometries similar to ours—by providing immediate feedback on the feasibility of selected impedance parameters.
To make this tool widely accessible, we provide an open-source Python implementation on GitHub [^5] https://github.com/mosesnah-shared/ImpedanceClassification/blob/main/ImpedanceClassification/success_failure_predictor.py. Users can input a set of impedance parameters, and the neural network will predict the probability of success, eliminating the need for time-consuming real-world testing. This is particularly valuable for inexperienced robot programmers, as it allows them to iteratively refine their parameter choices without direct physical experimentation.
To evaluate the neural network’s effectiveness, we randomly selected 30 predicted solutions and tested these parameter sets for each peg on the real robot. Remarkably, all parameter sets were successful, except for one case each for the square and triangle pegs. In total, these solutions resulted in 118 successful insertions out of 120 trials. A video of the experiments on the robot can be found on Dropbox [^6] https://www.dropbox.com/scl/fo/gav5l7kpzqs69dzbi06xj/AMGmfsqsbLDaThPr7xVikWo?rlkey=pc443f61levpdeme9exejhsa1&dl=0.
These results demonstrate that, despite the overall low success rate across the full parameter search space, the trained neural network significantly improved the success rate by guiding the selection of feasible impedance parameters. This reduces the need for exhaustive trial-and-error tuning, making impedance selection more systematic, efficient, and accessible for users.
IV Discussion
Our study investigated key factors for successful peg-in-hole assembly, emphasizing the role of impedance control in contact-rich robotic manipulation. We demonstrated that a range of successful parameters exists, underscoring the flexibility of solutions rather than relying solely on optimization. This complements traditional approaches by suggesting multiple viable solutions rather than a single optimal one.
While conventional robotic control strategies often prioritize higher stiffness values to improve precision and stability, our findings suggest a different perspective. Lower stiffness values, particularly for parameters such as $k_{x}$ (translation along $x$), $k_{y}$ (translation along $y$), $k_{A}$ (rotation about $z$), $\zeta_{t}$ (translational damping ratio), and $\zeta_{r}$ (rotational damping ratio), proved effective. Physically, lower translational impedance facilitates easier insertion of the peg into the hole, suggesting that higher stiffness may not always be necessary or beneficial for optimal performance in contact-rich tasks.
Consistent with previous studies [27, 62], our research indicates that increasing stiffness along unconstrained directions, such as $k_{z}$, can indeed improve success rates. However, for other parameters, moderate or even lower stiffness values were equally effective, reinforcing the idea that parameter selection should be driven by task-specific characteristics rather than a single, universally optimal approach.
Interestingly, our results suggest that the damping ratio’s influence on assembly success may be less critical than previously assumed. Thus, adopting default damping ratio values might be practical in many scenarios, simplifying parameter tuning without significantly compromising performance. However, more dynamic tasks with fewer physical constraints could necessitate more careful tuning of damping parameters.
V Conclusion
Our study explored key aspects of successful peg-in-hole assembly, highlighting the advantages of impedance control through Elementary Dynamic Actions (EDA) for contact-rich robotic manipulation.
We demonstrated that EDA supports multiple successful impedance parameter sets, indicating flexibility and adaptability in selecting effective control strategies. Principal Component Analysis (PCA) revealed a structured, lower-dimensional representation of these successful solutions, simplifying the complexity of the impedance parameter space.
Notably, our findings highlight that certain impedance settings generalize across different peg types, revealing common underlying strategies. By categorizing impedance parameters into intuitive groups—low, medium, and high—we facilitate efficient programming and rapid adaptation to variations in assembly tasks. This approach aligns with Bernstein’s [35] concept of “repetition without repetition,” mirroring the adaptability of human motor control.
All CAD data and software related to this study are publicly available, providing researchers with valuable resources to validate, extend, or apply these findings to real-world robotic assembly applications.
VI Limitations and future work
Our study also indicates that the effectiveness of EDA not only depends on the impedance parameter choice but also depends on the chosen strategy for kinematic primitives. While EDA’s modularity offers many benefits, the specific combination of submovements and oscillations used in our experiments played a crucial role in achieving successful outcomes. This dependency suggests that further refinement and customization of kinematic strategies may be necessary to optimize EDA’s performance across different tasks. Here, teaching by demonstration might be a promising way to facilitate the programming of kinematic primitives [31, 40].
Our study involved several parameter choices that could influence the outcomes of our investigation. For instance, during the Principal Component Analysis (PCA), we reduced the dimensionality of the solution space from eight to three principal components, capturing a certain percentage of the variance. This choice, while effective for our purposes, may not generalize across different contexts or datasets.
In the neural network model for binary classification, we used a structure with two hidden layers, each consisting of 64 neurons, and trained the network over $2^{13}$ epochs. The Adam optimizer with a fixed learning rate of $1\cdot 10^{-4}$ was used to update the weights. Although these are common choices in neural network design, other configurations and hyper-parameters could also have been effective in producing a feasible neural network.
Appendix A Categories and parameter ranges
Table A shows three different categories for each of the parameter ranges. The units for the translational stiffnesses and rotational stiffnesses are N/m and Nm/rad, respectively. The damping ratio is unitless. Category $k_{t}$ $k_{r}$ $\zeta$ Small $50\leq k_{t}\leq 300$ $5\leq k_{r}\leq 10$ $0.1\leq\zeta\leq 0.3$ Medium $301\leq k_{t}\leq 700$ $11\leq k_{r}\leq 80$ $0.4\leq\zeta\leq 0.7$ Large $701\leq k_{t}\leq 1000$ $81\leq k_{r}\leq 200$ $0.8\leq\zeta\leq 0.9$
Appendix B Principal Component Analysis
We conducted a pair-wise t-test between slopes of the variance-accounted-for in the successful trials vs. the random trials. If the slopes in the successful trials were significantly steeper (i.e., the probability that the slopes were the same was $\alpha\leq 0.05$), we concluded that the PCs extracted from the successful trials were not a random artifact. Specifically, we concluded that there existed a non-arbitrary impedance solution subspace. The results can be seen in Fig. 10.
Appendix C Low Dimensional Representation of Solutions
Figure 11 shows 2D-projections of the successful (green) and unsuccessful (red) trials of the cylinder, hex, square, and triangle pegs are projected into the solution space obtained using PCA (from 5). Note that often the failures fell outside the PCA-reduced solution space.
Figure 11: 2D-projections of the successful (green) and unsuccessful (red) trials for all pegs projected onto the solution space.
Appendix D Optimal Cluster Count Using Silhouette Analysis
The silhouette value $s_{i}$ measures how well a data point $i$ fits within its assigned cluster compared to other clusters. It is computed using the formula:
$s_{i}=\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{max(a_{i},b_{i})},$ (5)
where $a_{i}$ is the average distance between the $i-$th point and all other data points within the cluster, and $b_{i}$ is the avg distance between the $i-$th point and all other data points within the next nearest cluster. From the above definition it is clear that $-1\leq s_{i}\leq 1$. Values close to 1 indicate well-clustered data, negative values suggest misclustered data, and values near zero indicate the datum is near the border of two clusters [65, 66].
With this, the overall clustering effectiveness can be evaluated by the silhouette score, $SC$, defined as:
$SC=\max_{k}s(k),$ (6)
D-A Silhouette coefficients by peg type
When using k-means clustering to identify families of solutions in the cylinder, hex, square, and triangle pegs we found 2, 4, 3, and 2 solution clusters, respectively. These clusters had silhouette coefficients of 0.33, 0.41, 0.29, and 0.37 for the cylinder, hex, square, and triangle pegs, as can be seen in Fig. 12.
Appendix E MANOVA to identify distinct clusters
E-A MANOVA combined results
The MANOVA results revealed significant differences among the clusters on the combined dependent variables, suggesting that the clusters’ multivariate means occupied at least three dimensions:
  • Dimension 0: Wilks’ Lambda $=0.0315$, $\chi^{2}(24)=3693.09$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.2383$, $\chi^{2}(14)=1,531.56$, $p<0.001$
  • Dimension 2: Wilks’ Lambda $=0.9824$, $\chi^{2}(6)=18.91$, $p=0.0043$
E-B MANOVA results by peg type
Cylinder: The independent variable (cluster group) had two levels. MANOVA results showed that the multivariate means of the clusters occupied a one-dimensional space:
  • Wilks’ Lambda $=0.2198$, $\chi^{2}(8)=213.62$, $p<0.001$
Hex: The independent variable had four levels. The clusters spanned at least three dimensions:
  • Dimension 0: Wilks’ Lambda $=0.0197$, $\chi^{2}(24)=683.51$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.1854$, $\chi^{2}(14)=293.27$, $p<0.001$
  • Dimension 2: Wilks’ Lambda $=0.9087$, $\chi^{2}(6)=16.65$, $p=0.0106$
Square: The independent variable had three levels. The clusters spanned at least two dimensions:
  • Dimension 0: Wilks’ Lambda $=0.1029$, $\chi^{2}(16)=1454.05.39$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.4322$, $\chi^{2}(7)=536.42$, $p<0.001$
Triangle: The independent variable had two levels. MANOVA results showed a one-dimensional space:
  • Wilks’ Lambda $=0.2391$, $\chi^{2}(8)=135.91$, $p<0.001$
Appendix F Neural Network structure
For the neural network, we chose the following structure with two hidden layers:
The two hidden layers and output of the neural network were defined as follows: the first hidden layer transformed the eight-dimensional normalized input features $\mathbf{X}^{(i)}=\{k_{x}^{(i)},k_{y}^{(i)},k_{z}^{(i)},k_{C}^{(i)},k_{B}^{(i)},k_{A}^{(i)},\zeta_{t}^{(i)},\zeta_{r}^{(i)}\}\in\mathbb{R}_{>0}^{8}$ to 64 neurons; the second hidden layer took the 64 features from the hidden layer and passed them to another 64 neurons; the output layer transformed the 64 outputs of the second layer to the predicted output $\hat{y}^{(i)}$. After each of the first two layers, a Rectified Linear Unit (ReLU) activation function was used to introduce non-linearity into the model. A Sigmoid activation function was used at the output layer to map the final value to a probability between 0 and 1. The Adam optimizer [68] was used to update the weights of the network with a fixed learning rate of $1\times 10^{-4}$.
Appendix G Linear Model Fitting Performance for Impedance Parameters
Tab. II presents the complete details of the linear model results. Moreover, Tab. III reports the coefficients of determination and related p-values for the linear models used to fit the relative success rates of Fig. 4.